OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 2x - cos 2x A. 2 sin^2 x - 2 sin x cos x + 1 B. 2 sin x C. 2 sin^2 x + 2 sin x cos x - 1 D. 2 sin^2 x - 2 sin x cos x - 1

OpenStudy (mertsj):

The double angle identities are: \[\sin 2x=2\sin x \cos x\]

OpenStudy (mertsj):

\[\cos 2x=\cos ^2x-\sin ^2x\]

OpenStudy (mertsj):

\[\cos 2x=2\cos ^2x-1\]

OpenStudy (mertsj):

\[\cos 2x=1-\sin ^2x\]

OpenStudy (mertsj):

Use those to get one of your choices.

OpenStudy (anonymous):

or you could find the least common mulitiple of the problem and go from there, if you don't know how to do the other way

OpenStudy (anonymous):

last common multiple problem

OpenStudy (anonymous):

is the lcd 2x?

OpenStudy (anonymous):

you would find the LCM of sin2x and of cos2x

OpenStudy (anonymous):

is it A?

OpenStudy (anonymous):

@Mertsj ?

OpenStudy (mertsj):

No it is NOT A

OpenStudy (mertsj):